OpenStudy (anonymous):

Input the equation of the given line.The line including (2, 3) and (2, 5) Input the equation of the given line.The line including (2, 3) and (2, 5) @Mathematics

jimthompson5910 (jim_thompson5910):

The x coordinates are the same and equal to 2. So the equation of the line is x = 2, and this is a vertical line that cuts through 2 on the x axis

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

y=2x+

jimthompson5910 (jim_thompson5910):

no it's simply x = 2 and no y terms are involved

OpenStudy (anonymous):

Ok thanks can you help with another

OpenStudy (anonymous):

Input the equation of the given line in standard form.The line including (3, 1) and (-2, 3).

jimthompson5910 (jim_thompson5910):

step 1 is to find the slope, how do we do that?

OpenStudy (anonymous):

y1-y2 x1-x2

jimthompson5910 (jim_thompson5910):

good, so subtract the corresponding coordinates like you have what do you get when you do that?

OpenStudy (anonymous):

-2/5

jimthompson5910 (jim_thompson5910):

yep, that's what I got too

jimthompson5910 (jim_thompson5910):

so the slope is -2/5, which means that m = -2/5

jimthompson5910 (jim_thompson5910):

now plug this into y - y1 = m(x-x1)

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

y-1=-2/5(x-3)

OpenStudy (anonymous):

does y = 2x + 7

jimthompson5910 (jim_thompson5910):

y-1=-2/5(x-3) is correct

jimthompson5910 (jim_thompson5910):

now convert to standard form, are you familiar with that?

OpenStudy (anonymous):

No I dont think so

jimthompson5910 (jim_thompson5910):

The idea is to convert y - y1 = m(x - x1) into the form Ax + By = C which is known as standard form

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

so how do we convert?

OpenStudy (anonymous):

well I don't know.

jimthompson5910 (jim_thompson5910):

To convert, follow the steps below y-1=-2/5(x-3) 5y-5=-2(x-3) 5y-5=-2x+6 2x+5y = 6+5 2x+5y = 11

OpenStudy (anonymous):

Ok, then after that convert to y=mx=b

jimthompson5910 (jim_thompson5910):

no that's it because they want it in standard form

OpenStudy (anonymous):

Oh, Ok let me see if it is right

OpenStudy (anonymous):

Awesome man... Can you show me how to do this problem?

jimthompson5910 (jim_thompson5910):

you mean go over it again?

OpenStudy (anonymous):

No I have a new concept, and I am trying to study for a quiz. Using the given equations, find the point of intersection. If there is no point of intersection, type no points . 6 x - 2 y = 5 3 x - y = 10 The lines intersect at

jimthompson5910 (jim_thompson5910):

what do you notice about the ratio between corresponding coefficients?

OpenStudy (anonymous):

They are the same is that what you mean?

jimthompson5910 (jim_thompson5910):

I mean what's the relationship between 6 and 3 (the x coefficients)?

OpenStudy (anonymous):

Ok I think I got it. 6x-2y=5 -2(3x-y=10) 6x-2y=5 -6x+2y=-20

OpenStudy (anonymous):

Is that what we are looking for?

jimthompson5910 (jim_thompson5910):

yes now add the equations

OpenStudy (anonymous):

0+0=-15

jimthompson5910 (jim_thompson5910):

so you basically get 0 = -15, which is NEVER true

jimthompson5910 (jim_thompson5910):

so what does that mean?

OpenStudy (anonymous):

So it is a no point? (As stated in the question)

jimthompson5910 (jim_thompson5910):

yes, there are no points of intersection

OpenStudy (anonymous):

Let me see if it correct... Lol

OpenStudy (anonymous):

Yup, I will figure this out... can you follow to see if I am correct?

OpenStudy (anonymous):

Using the given equations, find the point of intersection. If there is no point of intersection, type no points . Leave answers in fractional form. 3 x + 2 y = 6 4 x + y = 1 The lines intersect at __________

jimthompson5910 (jim_thompson5910):

ok show me what you got

OpenStudy (anonymous):